Oxygen gas inside a 1.5 L gas tank has a pressure of 0.95 atm. Provide the temperature remains constant, how much pressure is needed to reduce it's volume by 1/2

1
Expert's answer
2022-05-23T17:02:09-0400

Given:

P1 = 0.95 atm

V1 = 1.5 L

T1 = T2 = constant

P2 = unknown

V2 = (1/2) × V1 = (1/2) × 1.5 L = 0.75 L


Formula: P1V1 = P2V2


Solution:

Since the temperature and amount of gas remain unchanged, Boyle's law can be used.

Boyle's gas law can be expressed as: P1V1 = P2V2

To find the pressure, solve the equation for P2:

P2 = P1V1 / V2

P2 = (0.95 atm × 1.5 L) / (0.75 L) = 1.90 atm

P2 = 1.90 atm


Answer: 1.90 atm